For F1 layer the maximum ionic density is 2.3 x 104 electrons per cc. The critical frequency for this layer will be


1
Expert's answer
2021-01-07T10:56:57-0500

Given,

Maximum ionic density "(N_{max})=2.3\times 10^{4} /cc"

Critical frequency "(fc)=9\sqrt{N_{max}}"

Now, substituting the values,

"f_c=9\sqrt{2.3\times 10^{4}\times 10^{6}}"

"\Rightarrow f_c=9\times 1.51\times 10^2\times 10^3" Hz

"\Rightarrow f_c=13.6\times 10^5Hz"

"\Rightarrow f_c=1.36MHz"